# Adhesion Energy of Phosphorene on different pristine and oxidized metallic substrates

### v1

@@ 4 Conclusion
DFT calculations were performed to determine the adhesion energy of both pristine and oxidized phosphorene on different pristine and oxidized metallic substrates. The results demonstrate that oxidized phosphorene generally exhibits higher adhesion to all substrates compared to its pristine counterpart. This enhanced adhesion can be attributed to more favorable interactions between the non-bonding states of oxygen atoms and the empty electronic states of the substrates.
The RMSD structural analysis revealed that oxidized phosphorene exhibits higher structural distortion values, consistent with its more favorable adhesion to substrates. Phosphorene adhesion is consistently more favorable on pristine metal surfaces than on their oxidized counterparts. This difference stems from the reduced availability of electronic charge for bonding in metal oxides, due to increased localization within the oxide structure.
Comparison of the calculated adhesion energies with experimental tribological results [vezzelli2024different] reveals that stronger adhesion (i.e., more negative energy values) correlates with reduced CoF. Consistently, the nearly negligible effect of BP on aluminum, copper, iron, and steel samples agrees with the low adhesion predicted for phosphorene on oxidized metal surfaces. In contrast, removing the oxide layer by scratching the iron sample prior to tribological testing resulted in a decrease in CoF, in line with the higher adhesion calculated for pristine iron compared to its oxidized form.
@@
The present computational work provides fundamental insights into the substrate-dependent tribological performance of BP by establishing quantitative relationships between interfacial adhesion energies and experimentally observed friction behavior, demonstrating that layer-substrate adhesion is a key factor governing the effectiveness of BP as a solid lubricant.

### v2

@@ 4 Conclusion
DFT calculations were performed to determine the adhesion energies of both pristine and oxidized phosphorene on different metal substrates, also considering the corresponding metal oxides. The results demonstrate that oxidized phosphorene generally exhibits higher adhesion to all substrates compared to its pristine counterpart. This enhanced adhesion can be attributed to more favorable interactions between the non-bonding states of oxygen atoms and the empty electronic states of the substrates.
The RMSD structural analysis revealed that oxidized phosphorene exhibits higher structural distortion, consistently with its more stronger adhesion to substrates. Phosphorene adhesion is always higher on pristine metal surfaces than on metal oxides. This difference stems from the reduced availability of electronic charge for bonding in metal oxides, due to increased localization within the oxide structure.
Comparison of the calculated adhesion energies with the lubricating performance of BP measured in the experiments described in ref. [vezzelli2024different] reveals that stronger adhesion correlates with reduced CoFs. Indeed the lowest CoF was measured for the iron substrate after the removal of the oxide layer by initial scratch of the sample, in line with the high adhesion calculated for pristine iron. Consistently, the nearly negligible effect of BP measured for aluminum, copper, iron, and steel samples agrees with the low adhesion predicted for phosphorene on oxidized metal surfaces.
@@
The present computational work provides fundamental insights into the substrate-dependent lubricating performance of BP by establishing quantitative relationships between interfacial adhesion energies and experimentally observed friction behavior, demonstrating that layer-substrate adhesion is a key factor governing the effectiveness of BP as a solid lubricant.